OpenStudy (anonymous):

im having the worst time with radicals I need sqrt 72x^2

OpenStudy (anonymous):

What are the factors of 72?

OpenStudy (anonymous):

2 2 2 3 3

OpenStudy (anonymous):

\[ \sqrt{72} = \sqrt{2\times 2 \times 2 \times 3 \times 3} = \sqrt{2\times 2^2\times 3^2} = 2\times 3\times \sqrt{2} = 6\sqrt{2} \]

OpenStudy (anonymous):

Break 72 into anything that you can take the square root of. Since 72 = 9*4*2, and x^2 is x*x, pull out all of the square roots you can from inside the radical (in this case x^2, 9 and 4). From there, the answer will be easy to break down. Hope this helps.

OpenStudy (anonymous):

On the other hand \[ \sqrt{x^2} = |x| \]

OpenStudy (anonymous):

Good point, wio.

OpenStudy (anonymous):

so you multiply the factors that show up twice and put it outside of the radical 6x sqrt 2?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

oh ok lol thank you

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."